# Bourbaki degree of pairs of projective surfaces

### v2

@@ 5 Pencils of cubics and degree 6 curves inside quadric surfaces
The results are derived from the previous sections, Section 3.1 and Section 4, and also from general results for reflexive sheaves of rank two on $\mathbb{P}^{3}$, found in the classical works [14], [21], [15], [3] and [16].
@@ 5.1 Pencils of cubics
In this section, we focus on the case of pencils of cubics, i.e., $d_{f}=d_{g}=2$. The Bourbaki degree of a sequence $\sigma$ in terms of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ is given by the formula
@@
The assumption $\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ enables us to apply the result [16, Theorem 1.1] to ${\mathcal{T}}_{\sigma}(1)$ and obtain the following upper bounds for $m(\sigma)$ and the third Chern class $c_{3}({\mathcal{T}}_{\sigma})$:
@@
    if $7\leq m(\sigma)\leq 8$, then $c_{3}\leq 16-2m(\sigma)$. In particular, when $m(\sigma)=8$, $\sigma$ is locally free.
@@
The following result is obtained using [15, Theorem 0.1] and the bound obtained in Section 3.1, $(a)$:
@@ 8 $e\leq 2$
Let $\sigma=(f,g)$ be a general pencil of cubics in $\mathbb{P}^{3}$. Then $m(\sigma)=0$, the number of singular members is $32$, and all singular members have one singular point, in particular $m(\sigma)=0$. Moreover $c_{3}({\mathcal{T}}_{\sigma})=32$. (see Example 11).
@@
This follows from intersection theory for the bundle of principal parts (see [10], Proposition $7.1$ and Proposition $7.4$). ∎
Rewriting Section 3.1, $(d)$ for pencils of cubics, we obtain:
@@
Let $\sigma$ be a normal pencil of cubics. Then $\sigma$ is compressible if and only if $m(\sigma)=12$ (see Example 34).
@@
Here, $m(\sigma)=12$, the annihilator ideal and the 0-th Fitting ideal coincide, with the support being three lines $V(x_{0},x_{3}),V(x_{1},x_{3})$ and $V(x_{0},x_{1})$, where the last one has multiplicity $10$ and the other two are simple.
Using Section 4, we obtain the following bounds for $\mu$-semistability of ${\mathcal{T}}_{\sigma}$ in terms of $m(\sigma)$:
@@
Since $\mu({\mathcal{T}}_{\sigma})=-2$, we show for $m(\sigma)\leq 6$ that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$. Since $m(\sigma)\leq 6$, we conclude that $\sigma$ is neither compressible nor free, from the previous results 33, 36 and 38. Let us suppose that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. Since $\sigma$ is not free, we can apply the result Section 4, $(a)$, and conclude $\operatorname{Bour}(\sigma)\leq 2$. On the other hand, we have
@@
if $m(\sigma)\leq 2$, we got a contradiction with the bound $\operatorname{Bour}(\sigma)\leq 5$ established earlier in Section 4, $(b)$. ∎
@@
The bound $m(\sigma)\leq 8$ is obtained for $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ in 29 for pencils of cubics, and $m(\sigma)\leq 9$ holds for $e\geq 1$, thus $m(\sigma)=9$ only if $e=1$. From the formula of the Bourbaki degree, we obtain $\operatorname{Bour}(\sigma)=0$ in this case, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$. On the other hand, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$, then $e=1$, and the equation $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=9$. ∎
Example 37 (Free, unstable incompressible pencil of cubics ($m(\sigma)=9$)).
@@
If $m(\sigma)=8$, by 31, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 2$. If $e=2$, we note that $c_{2}({\mathcal{T}}_{\sigma}(2))=0$ and ${\mathcal{T}}_{\sigma}(2)$ is strictly $\mu$-semistable with $c_{1}=0$, thus it follows from [3, Lemma 2.0, (a)] that ${\mathcal{T}}_{\sigma}(2)\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}$. If we assume $e=1$, we obtain $\operatorname{Bour}(\sigma)=1$ and, using Remark 7 for $B$ being a line (that is, $p_{a}(B)=0$ and $\deg(B)=1$), we obtain $c_{3}({\mathcal{T}}_{\sigma})=4$, on the other hand, 30 gives $c_{3}({\mathcal{T}}_{\sigma})=0$ for $m(\sigma)=8$, so we obtain a contradiction and this case cannot happen.
On the other hand, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=8$. ∎
@@
$\nabla\sigma=\begin{pmatrix}2x_{0}(x_{1}-x_{2})-x_{2}^{2}&x_{0}^{2}+x_{2}^{2}&-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3})&x_{2}^{2}\\ 0&-x_{2}x_{3}&x_{3}(2x_{2}-x_{1})&x_{2}(x_{2}-x_{1})\end{pmatrix}$
Using Macaulay2 software, we compute $e=2$ and $m(\sigma)=7$, so that $\operatorname{Bour}(\sigma)=1$.
@@
Let $\sigma=(f,g)$ be a normal pencil of cubics such that $m(\sigma)=7$. Then, the only possible $\mu$-semistable cases are:
  • $e=2$ and $\sigma$ is a nearly free sequence with $c_{3}({\mathcal{T}}_{\sigma})=2$ (see 40);
@@
For $e\geq 2$, since $c_{2}({\mathcal{T}}_{\sigma}(2))=1$ and $c_{1}({\mathcal{T}}_{\sigma}(2))=0$, from [3, Lemma 2.1], we conclude that the only possible cases are $c_{3}=0$ (stable case) or $c_{3}=2$ (strictly semistable case). Thus, for $e=2$, we must have $c_{3}({\mathcal{T}}_{\sigma})=2$, and thus we obtain the second case. For the third case, we must have a stable bundle ${\mathcal{T}}_{\sigma}(2)$ with Chern classes $(0,1,0)$, which are precisely null correlation bundles described in [22] fitting in a sequence of the form
@@
As we mentioned in Remark 27, the only possible case with $m(\sigma)=7$ for which we do not have examples is when $e=1$ and $c_{3}({\mathcal{T}}_{\sigma})=8$. We summarize the results of this subsection in the following theorem:
Theorem C. Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
@@
Item $(a)$ is 33 and item $(b)$ with 36 and Example 39. Item $(c)$ follows from 42 and item $(d)$ is 35. ∎
@@
The scheme structures $\operatorname{supp}(\mathcal{Q}_{\sigma})=\Xi_{\sigma}$ coincide, with support at a line $V(x_{2},x_{3})$ with multiplicity $4$. The sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)\xrightarrow{N}\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
where the first syzygy is $\nu=(-x_{1}^{2},x_{0}^{2},0,0)^{T}$, so we obtain $e=2,\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=4$. This is an example of a $3$-syzygy sequence, with $c_{3}({\mathcal{T}}_{\sigma})=16$.

### v3

@@ 4 Pencils of cubics and degree 6 curves inside quadric surfaces
The results are derived from the previous sections, Section 2.1 and Section 3, and also from general results for reflexive sheaves of rank two on $\mathbb{P}^{3}$, found in the classical works [undefn], [undefw], [undefo], [undefb] and [undefp].
Two major results which are used are the following ones, both assuming that $\kappa$ is an algebraically closed field of characteristic zero.
Theorem 28 ([undefp], Theorem 1.1).
Let $\mathcal{E}$ be a rank two reflexive sheaf on $\mathbb{P}^{3}$. Assume $c_{1}\geq-3$ and $h^{0}(\mathcal{E})=0$. Define the integers:
$A$ $\doteq\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil,\text{resp. ditto +1 if }c_{1}=1,3;$
$B$ $\doteq\left\lceil\frac{1}{3}(c_{1}^{2}+3c_{1}+8)\right\rceil,\text{resp. ditto +1 if }c_{1}=2,4;\text{ditto -1 if }c_{1}=-3.$
Then $c_{2}\geq A$. Furthermore:
  • (a)
    If $A\leq c_{2}\leq B$, then
    $c_{3}\leq(c_{1}+4)c_{2}-2\binom{c_{1}+3}{3}-2.$
  • (b)
    If $c_{2}>B$, then
    $c_{3}\leq c_{2}^{2}-c_{2}(2B-c_{1}-5)+B^{2}-B-2\binom{c_{1}+3}{3}-2.$
With this theorem, we obtain an upper bound for $m(\sigma)$ lower than $m_{0}$ when $\sigma$ is incompressible, so that $h^{0}({\mathcal{T}}_{\sigma})=0$, in the case $d_{f}=1,d_{g}=2$. In the case of pencils of cubics, however, we apply this to $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, so we must assume $\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ and treat the case $\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$ separately.
Another important result is the following, which enables us to obtain a lower bound for the existence of a syzygy in some cases:
Theorem 29 ([undefo], Theorem 0.1).
Let $\mathcal{E}$ be a rank two reflexive sheaf on $\mathbb{P}^{3}$ with $c_{1}=0$ of $c_{1}=-1$ and with $c_{2}\geq 0$. Let $t\in\mathbb{Z}$ such that either
  • (a)
    $c_{1}=0$ and $t>\sqrt{3c_{2}+1}-2$, or
  • (b)
    $c_{1}=-1$ and $t>\sqrt{3c_{2}+\frac{1}{4}}-\frac{3}{2}$.
Then $H^{0}(\mathcal{E}(t))\neq 0$.
@@ 4.1 Pencils of cubics
Now, assume $d_{f}=d_{g}=2$. The Bourbaki degree of a sequence $\sigma$ in terms of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ is given by the formula
@@
Applying Theorem 28 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, assuming $e>1$, we get
$c_{1}(\mathcal{E})=c_{1}({\mathcal{T}}_{\sigma}(1))=-d+2=-2$
and the quantities of the Theorem become
$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(4-4+3)\right\rceil=1$
$B$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+3c_{1}+8)\right\rceil=\left\lceil\frac{1}{3}(4-12+8)\right\rceil=0.$
Moreover, another important invariant is:
$c_{2}(\mathcal{E})$ $=c_{2}({\mathcal{T}}_{\sigma})-d+1$
$=m_{0}-m(\sigma)-d+1$
$=9-m(\sigma).$
Direct application of the result shows the next pair of propositions:
@@
    if $8\leq m(\sigma)\leq 9$, then $c_{3}\leq 16-2m(\sigma)$. In particular, when $m(\sigma)=8$, $\sigma$ is locally free.
@@
Next, we apply Theorem 29 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(2)$, a sheaf with $c_{1}=0$ and $c_{2}(\mathcal{E})=9-m(\sigma)\geq 0$ from the previous computations, assuming $e>1$. Then, the Theorem says that, for
$t$ $>\sqrt{3c_{2}+1}-2=\sqrt{27-3m(\sigma)+1}-2$
we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+2))\neq 0$.
Direct application of this bound for high values of $m(\sigma)=6,7,8$ gives minimum $t=2,1,1$ satisfying the inequality above, so we obtain
@@ 8 $e\leq 3$
Let $\sigma=(f,g)$ be a general pencil of cubics in $\mathbb{P}^{3}$. Then $m(\sigma)=0$, the number of singular members is $32$, and all singular members have one singular point, in particular $m(\sigma)=0$. Moreover $c_{3}({\mathcal{T}}_{\sigma})=32$ (see Example 10), and the minimal free resolution for the logarithmic sheaf ${\mathcal{T}}_{\sigma}$ is of the form
$0\to\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\to\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\to{\mathcal{T}}_{\sigma}\to 0,$
given by the Buchsbaum–Rim complex.
@@
This follows from intersection theory for the bundle of principal parts (see [undefi], Proposition $7.1$ and Proposition $7.4$). ∎
Rewriting Theorem A Theorem A, $(d)$ for pencils of cubics, we obtain:
@@
Let $\sigma$ be a normal pencil of cubics. Then $\sigma$ is compressible if and only if $m(\sigma)=12$ (see Example 35), and in this case ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)$.
@@
Here, $m(\sigma)=12$, with the irreducible components of $(\Xi_{\sigma})_{1}$ being three lines $V(x_{0},x_{3}),V(x_{1},x_{3})$ and $V(x_{0},x_{1})$, where the last one has multiplicity $10$ and the other two are simple.
Using Theorem B Theorem B, we obtain the following bounds for $\mu$-semistability of ${\mathcal{T}}_{\sigma}$ in terms of $m(\sigma)$:
@@
Since $\mu({\mathcal{T}}_{\sigma})=-2$, we show for $m(\sigma)\leq 6$ that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$. Since $m(\sigma)\leq 6$, we conclude that $\sigma$ is neither compressible nor free, from the previous results 34, 37 and 39. Let us suppose that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. Since $\sigma$ is not free, we can apply the result Theorem B Theorem B, $(a)$, and conclude $\operatorname{Bour}(\sigma)\leq 2$. On the other hand, we have
@@
if $m(\sigma)\leq 2$, we got a contradiction with the bound $\operatorname{Bour}(\sigma)\leq 5$ established earlier in Theorem B Theorem B, $(b)$. ∎
@@
The bound $m(\sigma)\leq 8$ is obtained for $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ in 30 for pencils of cubics, and $m(\sigma)\leq 9$ holds for $e\geq 1$, thus $m(\sigma)=9$ only if $e=1$. From the formula of the Bourbaki degree, we obtain $\operatorname{Bour}(\sigma)=0$ in this case, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$. On the other hand, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$, then $e=1$, and the equation $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=9$. ∎
Example 38 (Free, unstable and incompressible pencil of cubics).
@@
If $m(\sigma)=8$, by 32, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 3$. If $e=2$, we note that $c_{2}({\mathcal{T}}_{\sigma}(2))=0$ and ${\mathcal{T}}_{\sigma}(2)$ is strictly $\mu$-semistable with $c_{1}=0$, thus it follows from [undefb, Lemma 2.0, (a)] that ${\mathcal{T}}_{\sigma}(2)\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}$. If we assume $e\in{1,3}$, we obtain $\operatorname{Bour}(\sigma)=1$ and, using Remark 6 for $B$ being a line (that is, $p_{a}(B)=0$ and $\deg(B)=1$), we obtain:
$e=1$ $\Rightarrow c_{3}=4$
$e=3$ $\Rightarrow c_{3}=0$
On the other hand, 31 gives $c_{3}({\mathcal{T}}_{\sigma})=0$ for $m(\sigma)=8$, so we obtain a contradiction with the case $e=1$. For the case $e=3$, we would have a nearly-free locally free pencil, which cannot happen by 19.
Conversely, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=8$. ∎
@@
$\nabla\sigma=\begin{pmatrix}2x_{0}(x_{1}-x_{2})-x_{2}^{2}&x_{0}^{2}+x_{2}^{2}&-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3})&x_{2}^{2}\\ 0&-x_{2}x_{3}&x_{3}(2x_{2}-x_{1})&x_{2}(x_{2}-x_{1})\end{pmatrix}$.
There are four irreducible components in $(\Xi_{\sigma})_{1}$, all lines, corresponding to the four prime ideals below:
$\mathfrak{p}_{1}$ $=(x_{2},x_{3})$
$\mathfrak{p}_{2}$ $=(x_{1}-x_{2},x_{3})$
$\mathfrak{p}_{3}$ $=(x_{1},x_{2})$
$\mathfrak{p}_{4}$ $=(x_{0},x_{2})$
For the prime $\mathfrak{p}_{1}$, we may consider the element $u=(2x_{0}(x_{1}-x_{2})-x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{1}}$ and the elementary operation $R_{1}^{\prime}=uR_{1}$, to obtain the matrix
$(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}1&u(x_{0}^{2}+x_{2}^{2})&u(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))&ux_{2}^{2}\\ 0&-x_{2}x_{3}&x_{3}(2x_{2}-x_{1})&x_{2}(x_{2}-x_{1})\end{pmatrix}$,
which sends the first basis vector to $(1,0)\in R_{\mathfrak{p}_{1}}^{2}$ and therefore we may compute the cokernel by considering the second line of the remaining matrix, namely
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{(-x_{2}x_{3},x_{3}(2x_{2}-x_{1}),x_{2}(x_{2}-x_{1}))}$
$\simeq\frac{R_{\mathfrak{p}_{1}}}{(-x_{2}x_{3},x_{3},x_{2})}\simeq\frac{R_{\mathfrak{p}_{1}}}{\mathfrak{p}_{1}R_{\mathfrak{p}_{1}}}\simeq\kappa,$
so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
For the prime $\mathfrak{p}_{2}=(x_{1}-x_{2},x_{3})$, the same element $u\in R_{\mathfrak{p}_{2}}$ is invertible, so that we compute the cokernel analogously to obtain:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(-x_{2}x_{3},x_{3}(2x_{2}-x_{1}),x_{2}(x_{2}-x_{1}))}$
$\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3},x_{1}-x_{2})}\simeq\frac{R_{\mathfrak{p}_{2}}}{\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}}\simeq\kappa,$
so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
For the prime $\mathfrak{p}_{3}=(x_{1},x_{2})$, we take $u=(x_{0}^{2}+x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{3}}^{\times}$ to perform the elementary operations $R_{2}^{\prime}=R_{2}+u[x_{2}x_{3}]R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to obtain
$(\nabla\sigma)_{\mathfrak{p}_{3}}\sim\begin{pmatrix}u(2x_{0}(x_{1}-x_{2})-x_{2}^{2})&1&u(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))&ux_{2}^{2}\\ 0&0&x_{3}(2x_{2}-x_{1})-ux_{2}x_{3}(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))&x_{2}(x_{2}-x_{1})-ux_{2}^{3}x_{3}\end{pmatrix}$,
which sends $e_{2}\in R_{\mathfrak{p}_{3}}^{4}$ to $(1,0)\in R_{\mathfrak{p}_{3}}^{2}$. Thus, we may compute the cokernel by considering the remaining columns of the second row:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\frac{R_{\mathfrak{p}_{3}}}{(f_{1},f_{2})}$
where
$f_{1}$ $=x_{3}(2x_{2}-x_{1})+ux_{2}x_{3}(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))$
$f_{2}$ $=x_{2}(x_{2}-x_{1})+ux_{2}^{3}x_{3}.$
Considering $f_{1}\in I$, we obtain
$2x_{2}(1+ux_{2}(x_{1}-x_{0}+x_{3})-ux_{0}^{2})\equiv x_{1}\mod I.$
Substituting this into $f_{2}\in I$, we get:
$0\equiv x_{2}^{2}\left(1-(2(1+ux_{2}(x_{1}-x_{0}+x_{3}))-ux_{0}^{2})+ux_{2}x_{3}\right)\mod I,$
and the element multiplying $x_{2}^{2}$ is an unit in $R_{\mathfrak{p}_{3}}$, we conclude that $x_{2}^{2}\in I$. Hence $x_{1}^{2}\in I$ by $f_{1}\in I$, and we may rewrite $f_{1}$ and $f_{2}$ mod $I$ as:
$f_{1}$ $=x_{2}(2-ux_{0}^{2})-x_{1}\mod I$
$f_{2}$ $=-x_{1}x_{2}\mod I,$
and thus the ideal $I=(x_{1}^{2},x_{2}^{2},x_{1}x_{2},x_{2}v-x_{1})$ where $v\in R_{\mathfrak{p}_{3}}$ is a unit. Hence, one may write
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\kappa\langle 1,x_{1},x_{2}\rangle,$
so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=3$.
For the prime $\mathfrak{p}_{4}=(x_{0},x_{2})$, we take $u=(x_{3}(2x_{2}-x_{1}))^{-1}\in R_{\mathfrak{p}_{4}}$ and the elementary operations $R_{1}^{\prime}=R_{1}-u[-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3})]R_{2}\doteq R_{1}-uQR_{2}$ and $R_{2}^{\prime}=uR_{2}$, to obtain:
$(\nabla\sigma)_{\mathfrak{p}_{4}}\sim\begin{pmatrix}2x_{0}(x_{1}-x_{2})-x_{2}^{2}&(x_{0}^{2}+x_{2}^{2})+uQx_{2}x_{3}&0&x_{2}^{2}-uQx_{2}(x_{2}-x_{1})\\ 0&-ux_{2}x_{3}&1&ux_{2}(x_{2}-x_{1})\end{pmatrix}$,
sending $e_{3}\in R_{\mathfrak{p}_{4}}^{4}$ to $(0,1)\in R_{\mathfrak{p}_{4}}^{2}$. Hence, we compute the cokernel as:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}\simeq\frac{R_{\mathfrak{p}_{4}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{4}}}{I}$
with
$f_{1}$ $=2x_{0}x_{1}-x_{2}(2x_{0}+x_{2})=2x_{0}(x_{1}-x_{2})-x_{2}^{2}$
$f_{2}$ $=(x_{0}^{2}+x_{2}^{2})+uQx_{2}x_{3}$
$f_{3}$ $=x_{2}^{2}-uQx_{2}(x_{2}-x_{1})$
$u$ $=(x_{3}(2x_{2}-x_{1}))^{-1}$
$Q$ $=-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}).$
Thus,
$f_{2}$ $=x_{0}^{2}+x_{2}^{2}-ux_{3}x_{0}^{2}x_{2}+2u(x_{1}-x_{0}+x_{3})x_{3}x_{2}^{2}$
$=x_{0}^{2}(1-ux_{3}x_{2})+x_{2}^{2}(1+2u(x_{1}-x_{0}+x_{3})x_{3})$
$f_{3}$ $=x_{2}^{2}+u(x_{2}-x_{1})x_{0}^{2}x_{2}-2u(x_{2}-x_{1})(x_{1}-x_{0}+x_{3})x_{2}^{2},$
where the elements $(x_{2}-x_{1}),(x_{1}-x_{0}+x_{3}),x_{3}\in R_{\mathfrak{p}_{4}}^{\times}$. Since $f_{1}\in I$, we have
$x_{2}^{2}\equiv 2x_{0}(x_{1}-x_{2})\mod I,$
and thus
$f_{2}\equiv x_{0}(x_{0}(1-ux_{3}x_{2})+2(x_{1}-x_{2})+2u(x_{1}-x_{2})(x_{1}-x_{0}+x_{3})x_{3})\mod I,$
but since the element inside the parenthesis is invertible in $R_{\mathfrak{p}_{4}}$, we conclude that $x_{0}\in I$. From $f_{1}\in I$, we obtain that $x_{2}^{2}\in I$. Taking this into account, we conclude that
$f_{1}$ $\equiv-x_{2}^{2}\mod I$
$f_{2}$ $\equiv x_{2}^{2}\mod I$
$f_{3}$ $\equiv x_{2}^{2}(1-2ux_{2}(x_{1}+x_{3})+2ux_{1}(x_{1}+x_{3}))\equiv x_{2}^{2}\mod I,$
and therefore
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}\simeq\frac{R_{\mathfrak{p}_{4}}}{(x_{0},x_{2}^{2})}\simeq\kappa\langle 1,x_{2}\rangle,$
with $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}=2$. Using the associativity formula to compute $m(\sigma)$, we obtain:
$m(\sigma)=1\cdot 1+1\cdot 1+3\cdot 1+2\cdot 1=7.$
On the other hand, $e(\sigma)=2$, so using the Bourbaki formula we obtain $\operatorname{Bour}(\sigma)=1$, and conclude also the minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the form
$0\to\mathcal{O}_{\mathbb{P}^{3}}(-4)\to\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\to{\mathcal{T}}_{\sigma}\to 0.$
@@
Let $\sigma=(f,g)$ be a normal, $\mu$-semistable pencil of cubics such that $m(\sigma)=7$. Then, the only possible cases are:
  • $e=2$ and $\sigma$ is a nearly free sequence with $c_{3}({\mathcal{T}}_{\sigma})=2$ (see 41);
@@
From 32, we know that $e\leq 3$. For $\mu$-semistable sheaves, $e\geq 2$ and, since $c_{2}({\mathcal{T}}_{\sigma}(2))=1$ and $c_{1}({\mathcal{T}}_{\sigma}(2))=0$, from [undefb, Lemma 2.1], we conclude that the only possible cases are $c_{3}=2$ (strictly semistable case) or $c_{3}=0$ (stable case). For the strictly semistable case $e=2$ we obtain $\operatorname{Bour}(\sigma)=1$, so we get the first case.
For the second case, we must have a stable bundle ${\mathcal{T}}_{\sigma}(2)$ with Chern classes $(0,1,0)$, which are precisely null correlation bundles described in [undefx] fitting in a sequence of the form
@@
Theorem C.
Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
@@
Item $(a)$ is 34 and item $(b)$ with 37 and Example 40. Item $(c)$ follows from 43 and item $(d)$ is 36. ∎
@@
The scheme structures $\operatorname{supp}(\mathcal{Q}_{\sigma})=\Xi_{\sigma}$ coincide, with support at a line $V(x_{2},x_{3})$. Over the prime ideal $\mathfrak{p}=(x_{2},x_{3})$, we perform some elementary operations for the matrix $\nabla\sigma$, first $R_{1}^{\prime}=R_{1}-R_{2}$, and then $R_{2}^{\prime}=u/3\cdot R_{2}$, where $u=(x_{1}^{2})^{-1}\in R_{\mathfrak{p}}$, to obtain the form:
$(\nabla\sigma)_{\mathfrak{p}}\sim\begin{pmatrix}0&0&3x_{2}^{2}-x_{3}^{2}&3x_{3}^{2}-2x_{2}x_{3}\\ ux_{0}^{2}&1&(u/3)x_{3}^{2}&(2/3)ux_{2}x_{3}\end{pmatrix}$,
so it sends the second basis vector to $(0,1)\in R_{\mathfrak{p}}^{2}$, and then we may compute the cokernel using the first line and remaining columns:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}$ $\simeq\frac{R_{P}}{(3x_{2}^{2}-x_{3}^{2},3x_{3}^{2}-2x_{2}x_{3})}$
$\simeq\kappa\langle 1,x_{2},x_{3},x_{2}^{2}\rangle,$
and thus we conclude $\operatorname{{length}}(\mathcal{Q})_{\mathfrak{p}}=4$. Since $\deg(R/\mathfrak{p})=1$, using the associativity formula we obtain $m(\sigma)=4$. A first syzygy is $\nu=(-x_{1}^{2},x_{0}^{2},0,0)^{T}$, so we obtain $e=2$, and thus $\operatorname{Bour}(\sigma)=4$.
The sheaf ${\mathcal{T}}_{\sigma}$ admits a minimal free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)\to\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
obtained using Macaulay2, so this is an example of a $3$-syzygy sequence, with $c_{3}({\mathcal{T}}_{\sigma})=16$.
To finish this subsection, we explore in an example that the Bourbaki degree of a pencil of cubics does not only depend on the singularity type of each cubic, but rather how these align when we consider the associated pencil.
Example 45.
Consider the cubic polynomials:
$f$ $=x_{3}(x_{1}^{2}-x_{0}x_{2})+x_{1}^{2}(x_{0}-x_{1})$
$g$ $=x_{3}(x_{1}^{2}-x_{0}x_{2})+x_{1}^{3}$
$g^{\prime}$ $=x_{0}(x_{3}^{2}-x_{1}x_{2})+x_{3}^{3}.$
Here, $V(f)$ is a cubic surface with singularity type $4A_{1}$, with primary ideals $(x_{0},x_{1},x_{2})$, $(x_{0},x_{1}^{2},x_{3})$ and $(x_{1},x_{2},x_{3})$. On the other hand, the surfaces $V(g)$ and $V(g^{\prime})$ are both of type $A_{1}2A_{2}$ and are isomorphic, via the change of coordinates
$x_{0}$ $\mapsto x_{1}$
$x_{1}$ $\mapsto x_{3}$
$x_{3}$ $\mapsto x_{0}$
and fixing $x_{3}$. The sequence $\sigma=(f,g)$ is non-normal with $V(x_{1})$ in the divisorial component of the Jacobian scheme. On the other hand, $\sigma^{\prime}=(f,g^{\prime})$ is a normal, generic pencil with $m(\sigma^{\prime})=0$.